OpenStudy (anonymous):

Solve the Equation. Check your Answer. √3x+5=11

OpenStudy (anonymous):

maybe, x=12

OpenStudy (anonymous):

What is the radicand? is it 3, 3x, or 3x+5?

OpenStudy (anonymous):

I'm not sure ? I was really concfused on this problem

OpenStudy (anonymous):

Yeah it was right. (:

OpenStudy (anonymous):

which one does the problem look like: \[\sqrt{3} x+5=11\] \[\sqrt{3x}+5=11\] or \[\sqrt{3x+5}=11\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

ok \[\sqrt{3x}+5=11\] subtract 5 from both sides \[\sqrt{3x}=11-5\] \[\sqrt{3x}=6\] square both sides \[3x=6^{2}\] \[3x=36\] divide both sides by 3 \[x=12\]

OpenStudy (anonymous):

Thank you (: I understand now

OpenStudy (anonymous):

Glad to help :)